OpenStudy (mathmath333):

A bag containing 4 white and 5 black ball. A man draws three at a random. What are the odds against these being all black.What are the odds against these being all black.

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} & \normalsize \text{ From a bag containing 4 white and 5 black balls a man draws}\hspace{.33em}\\~\\ & \normalsize \text{ 3 at a random. What are the odds against these being }\hspace{.33em}\\~\\ & \normalsize \text{ all black ?}\hspace{.33em}\\~\\ & a.)\ \dfrac{5}{37} \hspace{.33em}\\~\\ & b.)\ \dfrac{37}{5} \hspace{.33em}\\~\\ & c.)\ \dfrac{11}{13} \hspace{.33em}\\~\\ & d.)\ \dfrac{13}{37} \hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (mathmath333):

Got it \(\large \color{black}{\begin{align} & \dfrac{1}{\dfrac{\dbinom{5}{3}}{\dbinom{9}{3}-\dbinom{5}{3}}} \hspace{.33em}\\~\\ \end{align}}\)